40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 21 18:18:42 2022, yehonathan, Update, BHD, POX Alignment POX11_alignment.png
    Reply  Thu Apr 28 17:45:53 2022, Anchal, Update, BHD, POX Beam issues PXL_20220429_002703686.jpg
       Reply  Sat Apr 30 13:29:33 2022, Paco, Update, BHD, POX Beam issues --- fixed, ASL alignment on vertex chamber 
          Reply  Mon May 2 17:51:30 2022, Anchal, Update, BHD, POX Beam aligned on RFPD, YARM locked 
Message ID: 16799     Entry time: Thu Apr 21 18:18:42 2022     Reply to this: 16818
Author: yehonathan 
Type: Update 
Category: BHD 
Subject: POX Alignment 

{Yehonathan, Paco}

BS, ITMX and ETMX were aligned to get flashing in the X arm.

I aligned the POX beam on the ITMX table using a mixture of the old POP and POX optics. The beam was stirred to the POX11 RFPD. We measure the DC power using a scope but we see nothing. We went and saw that the POX11 cable was not connected to RF rack so we connected it along with some other RFPD cables.

We return but there is still no DC. We ndscope C1:LSC-POX11_I_ERR_DQ C1:LSC-POX11_Q_ERR_DQ and maximize the signal (attachment). The readout is very weak though. It should be as strong as POY which we already observed to have good SNR.

We also noticed that the one of the beam dumps for the POX RFPD is not glued and easily falls down.

Attachment 1: POX11_alignment.png  112 kB  Uploaded Thu Apr 21 19:27:21 2022  | Hide | Hide all
POX11_alignment.png
ELOG V3.1.3-